OpenStudy (anonymous):

Estimate the length of the radius, if the area of a circle is 750 mm2. A. 19.78 mm B. 23.87 mm C. 9.36 mm D. 15.45 mm

hartnn (hartnn):

use the formula \(\Large A = \pi r^2\) here A = 750 find r

OpenStudy (anonymous):

would R be 375

hartnn (hartnn):

no... 750 = pi r^2 start by dividing by pi on both sides , what u get ?

OpenStudy (anonymous):

what is the radius

hartnn (hartnn):

that you need to find by isolating the r

hartnn (hartnn):

r^2 = 750/pi take square root on both sides

OpenStudy (anonymous):

ohh so would the anwser be 19.78

hartnn (hartnn):

no..how u got that ?

hartnn (hartnn):

750/pi is 238.73

hartnn (hartnn):

and its square root is ?

OpenStudy (anonymous):

ohh 15.45

hartnn (hartnn):

yes :)